OpenStudy (anonymous):

At the beach cleanup, Maya's group collected 14.8 pounds of trash on Saturday and 12.5 pounds on Sunday. Will's group collected the same total amount of trash as Maya's group but collected only 11.8 pounds on Saturday. How much trash did Will's group collect on Sunday?

OpenStudy (anonymous):

first you take 14.8 since thats what maya collected on saturday and take what will collected on saturday which is 11.8 and minus it which is 3 then you add 3 to 12.5 which is what maya got on sunday and your answer will be 15.5

OpenStudy (anonymous):

maya total = 14.8+12.5 = ? will = total same as maya = ? of this 11.8 was on Sat. , the remaining on Sunday. = ?

OpenStudy (anonymous):

in math, \[14.8+12.5=11.8+x\] solve for x

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem any time :)